OpenStudy (xxstudyingstuffxx):

Fan and Medal! The height of a cone shaped container is 70 centimeters and its radius is 15 centimeters. Cathy fills the container completely with honey. Every day Cathy uses 500 cubic centimeters of honey from the container to make honey cookies. After how many days will the container be completely empty? Round your answer to the nearest whole number. (Use π = 3.14.) 17 20 33 40

OpenStudy (xxstudyingstuffxx):

@Qwertty123

OpenStudy (qwertty123):

okay so for a Right Circular Cylinder. \[V=\pi r^2h\]

OpenStudy (xxstudyingstuffxx):

Ok. BTW: I know nothing about this. I am asking it on behalf of a friend.

OpenStudy (qwertty123):

oh! okay. XD well then 1.34* the radius which is... 15

OpenStudy (xxstudyingstuffxx):

But, would you be willing to point me into the direction of a trustworthy website where I can learn about it?

OpenStudy (anonymous):

\[\frac{\pi 15^2 70}{3\ 500}=\frac{21 \pi }{2}=33 \]

OpenStudy (xxstudyingstuffxx):

I dont know what that means. What is this symbol? π